OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation 4x - 3y = 24.

OpenStudy (anonymous):

x intercept = (6, 0); y intercept = (0, -8) x intercept = (-6, 0); y intercept = (0, 8) x intercept = (-8, 0); y intercept = (0, -6) x intercept = (8, 0); y intercept = (0, 6)

OpenStudy (anonymous):

x= (6,0) y= (0,-8) that should be right, you substitute the each variable with 0. The person above me is incorrect, I checked. Only "x intercept = (6, 0); y intercept = (0, -8)" is right.